If the random variable X follows Poisson distribution with
p(x = 0) =p(x =1),
Then find E(x) and p(x>1)?

Answer by Boreal(15235)

if the parameter is 1 e^(-1)*0^1/0!=e^(-1)*1^1/1!
Therefore the Poisson distribution of concern is where lambda=1
E(x)=mean=1
variance is also 1, along with sd
p(0)=p(1)=1/e=0.3679
so p(x>1)=1-2(0.36788)=0.2642